# Univ. Lille, Inria, CNRS, Centrale Lille, UMR 9189 CRIStAL, F-59000 Lille, France

### v1

@@ Abstract
To address this issue, we consider concolic unit testing, a combination of concrete execution with SMT-based symbolic execution, and propose CUTECat, a concolic execution tool targeting implementations of computational laws. Such laws typically follow a pattern where a base case is later refined by many exceptions in following law articles, a pattern that can be formally modeled using default logic. We show how to handle default logic inside a concolic execution tool, and implement our approach in the context of Catala, a recent domain-specific language tailored to implement computational laws. We evaluate CUTECat on several programs, including the Catala implementation of the French housing benefits and Section 132 of the US tax code. We show that CUTECat can successfully generate hundreds of thousands of testcases covering all branches of these bodies of law. Through several heuristics, we improve CUTECat’s scalability and usability, making the testcases understandable by lawyers and programmers alike. We believe CUTECat thus paves the way for the use of formal methods during legislative processes.
@@ 1 Introduction
One of the core issues lies with the particular structure of computational law, commonly consisting of a base case refined by exceptions spread out throughout law texts. This structure, corresponding to default logic [7, 26], is not straightforward to encode using modern programming languages. This leads to discrepancies between code and law, where design choices made by programmers can be hard to understand by lawyers. Conversely, maintaining such implementations when legislative processes modify the law becomes tricky, as programmers cannot necessarily accurately pinpoint required changes in their code based on changes in the legal texts.
@@
Our contributions are the following: we show how to concolically execute computational law programs by providing a formal concolic semantics for default terms (Sec. 2.2). Building on this semantics, we then propose several semantic transformations to improve the efficiency of concolic evaluation (Sec. 2.3). Relying on this formal model, we then implement CUTECat, a concolic execution execution engine for Catala programs (Sec. 3). CUTECat aims to detect all runtime errors, such as divisions by zero but also law inconsistencies due to missing, or conflicting interpretations. CUTECat includes several optimizations, both to improve its performance and scalability, as well as its usability, aiming to generate human-friendly testcases to facilitate legal interpretation by lawyers (Sec. 4). Finally, we experimentally evaluate CUTECat on a range of Catala programs (Sec. 5). Relying on real-world codebases from French and US laws, we first perform an ablation study to evaluate the impact of our different optimizations. We then conclude by empirically demonstrating that CUTECat can scale to the largest real-world Catala codebase currently available, namely, the implementation of the French housing benefits (19,655 lines of Catala, including specification), generating 186,390 tests in less than 7 hours of CPU time.
All of our development is open-source and publicly available on GitHub. Additionally, all experimental claims made in this paper will be documented in an artifact, which we plan to submit as part of the artifact evaluation process.
@@ 2.1 Computational Law and Default Logic
To address several societal issues, this computation can however be modified depending on households. For instance, low-income households (which we define in this example as earning no more than $10,000) might be taxed less, leading to a reduced tax rate of 10%. Furthermore, large households (which we define in this example as households with three or more children), might benefit from a tax break, for instance, lowering their tax rate to 15%. For now, we do not set any interpretation priority for households having both a low income and a large number of children.
@@
If more than one exception expression reduces to a non-empty value, that is if several exceptions are raised at the same time, then the default expression reduces to a conflict error (DefaultExceptionsConflict). Thus, yields .
@@ 2.2 Concolically Executing Default Terms
We now present how to support default terms during concolic execution. We first start with some preliminaries about concolic execution, and the related concept of symbolic execution.
@@
[C-DefaultExpr] e v ≠
@@
To illustrate how our semantics operates, we will rely on the following example, where and are respectively integer and boolean variables: . Let us assume that we want to concolically execute this term with and . Initially, both the global constraints and local constraints are trivial.
Following rule C-DefaultExpr, we must first reduce the exception to a value. To do so, since we have a default term (cf. Remark 1) and its default condition is , we can apply rule C-DefaultTrueNoExceptions, obtaining the concolic term . We conclude the application of rule C-DefaultExpr, leading to the concolic term . We then perform a similar reduction on the second exception by combining rules C-DefaultExpr and C-DefaultFalseNoExceptions to obtain the concolic term . As exactly one exception reduced to a non-empty or error value, the final step in the concolic execution is then to apply rule C-DefaultOneException; local constraints are thus appended to the current (empty) path condition. The execution therefore terminates with the value , accurately capturing that any set of inputs satisfying the path constraint will follow the same execution path.
@@
2.3 Optimizing Concolic Execution of Default Terms
The approach presented in the previous section allows us to perform concolic execution on default terms. However, to improve interactivity with developers, our goal is not only to exhaustively analyze all program execution paths, but also to find possible errors as early as possible during the analysis. In this section, we therefore propose several optimizations aiming to prune constraints leading to redundant cases, and to prioritize executions that might lead to conflict errors.
Lazily Evaluating Exceptions.
Consider for instance a default term , where and both evaluate to values that are neither empty nor a conflict, respectively generating constraints and . By following rule C-DefaultExpr, concolically executing would require concolically executing and generating its associated constraint , which would then be added to the path condition when applying rule C-DefaultExprConflict. Further iterations of concolic execution would therefore negate the last constraints in the path condition, i.e., those in , while preserving and . As any input satisfying leads to and evaluating to non-empty nor conflict values, thus raising a conflict, this approach would induce a number of redundant iterations that is exponential in the number of constraints in . By instead adopting rule DefaultExceptionsConflict-Lazy and stopping the execution after the evaluation of , we therefore prune from the path condition, thus reducing the number of iterations needed.
[C-DefaultExpr-Lazy] e v ≠
[C-DefaultExprOne-Lazy] e v_i ≠,
[C-DefaultExceptionsConflict-Lazy] e v ≠, v_i ≠,
Reorganizing Exceptions.
To leverage this fact, one key observation is that the result of the evaluation of a default term is independent of the evaluation order of the exception list. We formalize this property in Th. 2.1, which states that swapping any two exceptions results in the evaluating the default term to the same value, therefore allowing to use any exception permutation as a semantic-preserving transformation.
Theorem 2.1 (Independence of the exception evaluation order)
If there exists a default value such that , then
This therefore allows to use different evaluation orders depending on the program and current symbolic state. In our implementation, which we describe in more detail in the next sections, we permute exceptions to group them depending on their free variables, a technique reminiscent of constraint reordering introduced by Cadar et al. [9].
Pruning Non-Conflict Path Conditions.
This rule directly derives from Th. 2.1; indeed, the rule C-DefaultError-Early is conceptually equivalent to reorganizing the exception list to place at the head of the list, and then applying C-DefaultError. This is where the need for differentiating between the path condition and local constraint arises: while constraints corresponding to the evaluation of earlier exceptions can be safely dropped, the path condition up to this default term must be preserved.
[C-DefaultError-Early] e
@@ 3 CUTECat: Implementing Concolic Execution
[numbers=left,xleftmargin=2em,xrightmargin=0em,autogobble,firstline=2,lastline=51,firstnumber=1]catala_enrunning_example.catala_en
@@
By default, CUTECat outputs these different cases directly to the user, however, we also provide facilities to directly generate Catala test scopes. These test scopes can be saved, and easily replayed using the Catala toolchain, either by using the reference interpreter or by compiling it to one of the Catala backends.
As identified by our concolic execution, our running example can raise conflicts, i.e., when considering large, low-income households. This conflict is not due to an implementation error, but rather to an ambiguity in our (simplified) law. Resolving this ambiguity requires legal interpretation, which must be performed by lawyers, and not programmers. In this case, legal precedents might suggest that ambiguities shall be resolved in the most favorable way to citizens, therefore defining a 10% tax rate. To implement this decision, programmers could thus modify the Catala implementation of Article 3 as shown in Fig. 10, by materializing a priority order on the evaluation of both exceptions: as the exception is now defined on income_tax, this exception will trigger first, before reaching the base computation depending on tax_rate.
[numbers=left,xleftmargin=2em,xrightmargin=0em,autogobble,firstline=2,lastline=5,firstnumber=1]catala_enrunning_fixed.catala_en
CUTECat is implemented in 3,400 lines of OCaml code, and relies on the Z3 SMT solver [41], leveraged through its OCaml bindings. Additionally, it is an integral part of the Catala toolchain, and therefore leverages its build system, simplifying the use of CUTECat on Catala projects. To provide users with more control on the concolic execution, CUTECat allows the use of assertions in the source code to restrict the input space, for instance, to specify that only households in a given country or state must be considered. Additionally, users can provide hints about initial inputs for the concolic execution, in order to quickly explore variants of a given situation.
@@
Divisions by zero are a well-known source of runtime errors. To reason about them, concolic tools therefore consider divisions as implicitly branching, and collect whether the denominator is equal to zero as a symbolic constraint.
Unfortunately, raising an error when dividing by zero conflicts with the default terms optimizations presented in Sec. 2.3. When reorganizing exceptions, a default term that was previously evaluating to a conflict might now raise a division by zero, therefore contradicting Th. 2.1. One key observation is that the conflict value acts as a de facto uncatchable exception: when it appears, it stops the evaluation of the program, which immediately returns. We therefore conflate divisions by zero and conflicts, and return in both cases. From a user perspective, the main concern is whether the program can lead to an error, independently of which one; as we generate concrete inputs, it is always possible to run the reference Catala interpreter to provide precise error reporting.
@@
While CUTECat currently supports a large subset of Catala which is sufficient to analyze real-world programs (Sec. 5), operations on lists and dates are only partially supported. List operations in Catala include filtering depending on a condition, aggregating over list contents to compute all sorts of values, and applying a function to all elements in the list; providing symbolic encoding of such operations would require higher-order symbolic reasoning [55, 42]. On the other hand, dates can easily be represented in an SMT solver as the number of days since a specific point in time, e.g., the Unix epoch. This representation allows to model several operations, such as the addition of a (possibly symbolic) number of days, or the duration corresponding to the difference between two dates. Unfortunately, other operations, including month and year addition or returning the first day of a week or a month, have a more complex semantics [39] which does not naturally fit into this SMT encoding.
@@ 4.1 Generating Human-compatible Testcases
To do so, we rely on the encoding of optional constraints in the solver, dubbed “soft constraints” [59]. Z3 natively allows to combine SMT reasoning with solving optimization objectives, including the specification of soft constraints; however, we encountered significant slowdowns when attempting to use this feature, raising scalability issues. We posit that this is due to non-linearity in our soft-constraints – e.g., when specifying – and in our rounding function defined in Sec. 3. We instead adopt an alternative approach, by implementing custom soft constraints through multiple queries to the solver. Concretely, when an iteration of concolic execution terminates and a new path condition is generated, we first query Z3 to check whether is satisfiable. If so, we further query Z3 by iteratively refining with additional constraints forcing human-friendlier input generation.
@@ 4.2 Pattern-Matching Case Folding
Consider for example the snippet in Fig. 11, extracted from the Catala implementation of the French housing benefits. At lines 4-12, this program performs a match on the 9 variants of a sum type representing geographic areas of France (mainland, or overseas territories). To analyze this program, a naive implementation of our concolic interpreter would therefore consider 9 different, disjoint branches, leading to 9 distinct iterations.
@@
[numbers=left,xleftmargin=2em,xrightmargin=0em,autogobble,firstline=2,lastline=14,firstnumber=1]catala_enpattern_matching_folding.catala_en
@@ 4.3 SMT-Solving Optimizations
Z3 provides incremental solving capabilities, where sets of constraints can be incrementally added or removed through stack-based push and pop operators. When popping the stack, the solver backtracks to an earlier proof state, only removing lemmas learned between push and pop and avoiding reproving many facts.
We implemented support for incremental solving, pushing a new stack for each element in the path condition. As CUTECat’s exploration strategy is currently fixed to depth-first search (DFS), incremental solving is particularly appealing: constraints are naturally changed in a last-in-first-out fashion, corresponding to a stack-based data structure.
4.4 Default Term Concolic Execution
We presented in Sec. 2.3 optimizations to the concolic execution of default terms. CUTECat implements the lazy evaluation of exceptions, as well as exception reorganization to order exceptions according to their sets of free variables.
Pruning non-conflict path conditions is trickier to implement while following a DFS exploration strategy. A key invariant of DFS exploration is that, during a concolic iteration, the previous path condition (used to generate the current input) is a prefix of the current path condition up to the negated constraint. This does not hold with rule C-DefaultError-Early, as the local constraint is dropped from the path condition. We leave the study of alternative exploration strategies and their impact on incremental SMT solving to future work.

### v2

@@ Abstract
To address this issue, we consider concolic unit testing, a combination of concrete execution with SMT-based symbolic execution, and propose CUTECat, a concolic execution tool targeting implementations of computational laws. Such laws typically follow a pattern where a base case is later refined by many exceptions in following law articles, a pattern that can be formally modeled using default logic. We show how to handle default logic inside a concolic execution tool, and implement our approach in the context of Catala, a recent domain-specific language tailored to implement computational laws. We evaluate CUTECat on several programs, including the Catala implementation of the French housing benefits and Section 132 of the US tax code. We show that CUTECat can successfully generate hundreds of thousands of testcases covering all branches of these bodies of law. Through several heuristics, we improve CUTECat’s scalability and usability, making the testcases understandable by lawyers and programmers alike. We believe CUTECat paves the way for the use of formal methods during legislative processes.
@@ 1 Introduction
One of the core issues lies with the particular structure of computational law, commonly consisting of a base case refined by exceptions spread out throughout law texts. This structure, corresponding to default logic [6, 28], is not straightforward to encode using modern programming languages. This leads to discrepancies between code and law, where design choices made by programmers can be hard to understand by lawyers. Conversely, maintaining such implementations when legislative processes modify the law becomes tricky, as programmers cannot necessarily accurately pinpoint required changes in their code.
@@
Our contributions are the following: we show how to concolically execute computational law programs by providing a formal concolic semantics for default terms (Sec. 2.2). Relying on this formal model, we then implement CUTECat, a concolic execution execution engine for Catala programs (Sec. 3). CUTECat aims to detect all runtime errors, such as divisions by zero but also law inconsistencies due to missing, or conflicting interpretations. CUTECat includes several optimizations, both to improve its performance and scalability, as well as its usability, aiming to generate human-friendly testcases to facilitate legal interpretation by lawyers (Sec. 4). Finally, we experimentally evaluate CUTECat on a range of Catala programs (Sec. 5). Relying on real-world codebases from French and US laws, we first perform an ablation study to evaluate the impact of our different optimizations. We then conclude by empirically demonstrating that CUTECat can scale to the largest real-world Catala codebase currently available, namely, the implementation of the French housing benefits (19655 lines of Catala, including specification), generating 186390 tests in less than 7 hours of CPU time. Our development is open-source and publicly available on GitHub; all experimental claims made in this paper are documented in an artifact [23].
@@ 2.1 Computational Law and Default Logic
To address several societal issues, this computation can however be modified depending on households. For instance, low-income households (which we define in this example as earning no more than $10,000) might be taxed to a reduced rate of 10%. Furthermore, large households (which we define in this example as households with three or more children), might benefit from a tax break, for instance, lowering their tax rate to 15%. For now, we do not set any interpretation priority for households having both a low income and a large number of children.
@@
If more than one exception expression reduces to a non-empty value, that is if several exceptions are raised at the same time, then the default expression reduces to a conflict error (DefaultExceptionsConflict)111Two raised exceptions are considered a conflict even if they lead to the same result.. Thus, yields .
@@ 2.2 Concolically Executing Default Terms
We now present how to support default terms during concolic execution. We start with some preliminaries about concolic execution, and the related concept of symbolic execution.
@@
[C-DefaultExpr] e v ≠ inline]PG I added v1 in C-DefaultExpr to be consistent
@@
To illustrate how our semantics operates, we will rely on the following example, where and are respectively integer and boolean variables: . Let us assume that we want to concolically execute this term with and . Initially, both the global constraints and local constraints are trivial. Following rule C-DefaultExpr, we must first reduce the exception to a value. Since we have a default term (cf. Remark 1) and its default condition holds, we can apply rule C-DefaultTrueNoExceptions, obtaining the concolic term . We conclude the application of rule C-DefaultExpr, leading to the concolic term . We then perform a similar reduction on the second exception by combining rules C-DefaultExpr and C-DefaultFalseNoExceptions to obtain the concolic term . As exactly one exception reduced to a non-empty or error value, the final step in the concolic execution is then to apply rule C-DefaultOneException; local constraints are thus appended to the current (empty) path condition. The execution therefore terminates with the value , accurately capturing that any set of inputs satisfying the path constraint will follow the same execution path.
color=Plum,]AF: I removed mentions of soundness/completeness, they feel more confusing at this stage, and they are hinted at with the sentence "accurately capturing … same execution path"
@@ 3 CUTECat: Implementing Concolic Execution
[numbers=left,xleftmargin=2em,xrightmargin=0em,autogobble,firstline=2,lastline=51,firstnumber=1]catala_encode/running_example.catala_en
@@
By default, CUTECat outputs these different cases directly to the user, however, we also provide facilities to directly generate Catala test scopes. These test scopes can be saved, and easily replayed using the Catala toolchain, either by using the reference interpreter or by compiling color=Plum,]PG try to compile generated testcases? For now I interpret to see my overhead, but the execution time of compiled testcases may also be a good metric? it to one of the Catala backends.
As identified by our concolic execution, our running example can raise conflicts, i.e., when considering large, low-income households. This conflict is not due to an implementation error, but rather to an ambiguity in our (simplified) law. Resolving this ambiguity requires legal interpretation, which must be performed by lawyers, and not programmers. In this case, legal precedents might suggest that ambiguities shall be resolved in the most favorable way to citizens, therefore defining a 10% tax rate. To implement this decision, programmers could thus modify the Catala implementation of Article 3 as shown in Fig. 8, by materializing a priority order on the evaluation of both exceptions: as the exception is now defined on income_tax, this exception will trigger first, before reaching the base computation depending on tax_rate222Note that a more idiomatic fix would make use of Catala ”labels”, that enable programmers to order exceptions explicitly. We omit this solution to keep our presentation of Catala small. The Catala tutorial [36] has more information on labels..
[numbers=left,xleftmargin=2em,xrightmargin=0em,autogobble,firstline=2,lastline=5,firstnumber=1]catala_encode/running_fixed.catala_en
CUTECat is implemented in 3,400 lines of OCaml code, and relies on the Z3 SMT solver [18], called through its OCaml bindings. It is an integral part of the Catala toolchain, and therefore leverages its build system, simplifying the use of CUTECat on Catala projects. To provide users with more control on the concolic execution, CUTECat allows the use of assertions in the source code to restrict the input space, for instance, to specify that only households in a given country or state must be considered. Users can provide hints about initial inputs for the concolic execution, in order to quickly explore variants of a given situation. color=Plum,]PG implement that? R: It can just be CLI for default values per type I think
@@
Divisions by zero are a well-known source of runtime errors. To reason about them, concolic tools therefore consider divisions as implicitly branching, and collect whether the denominator is equal to zero as a symbolic constraint.333 Note that our semantics conflates runtime errors with conflicts, and returns in both cases. Indeed, the conflict value acts as a de facto uncatchable exception: when it appears, it stops the evaluation of the program, which immediately returns. From the user’s perspective, the main concern is whether the program can lead to an error, independently of which one. As we generate concrete inputs, it is always possible to run the reference Catala interpreter to provide precise error reporting.
@@
CUTECat currently supports a large subset of Catala which is sufficient to analyze real-world programs (Sec. 5) that mainly rely on integer-rational reasoning as well as algebraic datatypes. However, operations on lists and dates are only partially supported. List operations in Catala include filtering depending on a condition, aggregating over list contents to compute all sorts of values, and applying a function to all elements in the list; providing symbolic encoding of such operations would require higher-order symbolic reasoning [55, 43]. On the other hand, dates can easily be represented in an SMT solver as the number of days since a specific point in time, e.g., the Unix epoch. This representation allows to model several operations, such as the addition of a (possibly symbolic) number of days, or the duration corresponding to the difference between two dates. Unfortunately, other operations, including month and year addition or returning the first day of a week or a month, have a more complex semantics [41] which does not naturally fit into this SMT encoding.
@@ 4 Improving the Scalability and Usability of CUTECat
4.1 Optimizing Concolic Execution of Default Terms
The approach presented in Sec. 2.2 allows us to perform concolic execution on default terms. However, to improve interactivity with developers, our goal is not only to exhaustively analyze all program execution paths, but also to find possible errors as early as possible during the analysis. In this section, we therefore propose several optimizations aiming to prune constraints leading to redundant cases, and to prioritize executions that might lead to conflict errors.
Lazily Evaluating Exceptions.
Consider for instance a default term , where and both evaluate to values that are neither empty nor a conflict, respectively generating constraints and . By following rule C-DefaultExpr, concolically executing would require concolically executing and generating its associated constraint , which would then be added to the path condition when applying rule C-DefaultExceptionsConflict. Further iterations of concolic execution would therefore negate the last constraints in the path condition, i.e., those in , while preserving and . As any input satisfying leads to and evaluating to non-empty nor conflict values, thus raising a conflict, this approach would induce a number of redundant iterations that is exponential in the number of constraints in . By instead adopting rule C-DefaultExceptionsConflict-Lazy and stopping the execution after the evaluation of , we therefore prune from the path condition, thus reducing the number of iterations needed.
[C-DefaultExpr-Lazy] e v ≠
[C-DefaultExprOne-Lazy] e v_i ≠,
[C-DefaultExceptionsConflict-Lazy] e v ≠, v_i ≠,
Reorganizing Exceptions.
To leverage this fact, one key observation is that the result of the evaluation of a default term is independent of the evaluation order of the exception list. We formalize this property in Th. 4.1, which states that swapping any two exceptions results in evaluating the default term to the same value. This therefore allows to use different evaluation orders depending on the program and current symbolic state. In our implementation, which we describe in more detail in the next sections, we permute exceptions to group them depending on their free variables, a technique reminiscent of constraint reordering introduced by Cadar et al. [8].
Theorem 4.1 (Independence of the exception evaluation order)
If there exists a default value such that , then
color=Plum,inline]PG proof sketch or lemma? If we have room…
Pruning Non-Conflict Path Conditions.
This rule directly derives from Th. 4.1; indeed, the rule C-DefaultError-Early is conceptually equivalent to reorganizing the exception list to place at the head of the list, and then applying C-DefaultError. This is where the need for differentiating between the path condition and local constraint arises: while constraints corresponding to the evaluation of earlier exceptions can be safely dropped, the path condition up to this default term must be preserved.
[C-DefaultError-Early] e
Implementation.
CUTECat implements the lazy evaluation of exceptions, as well as exception reorganization to order exceptions according to their sets of free variables. Pruning non-conflict path conditions is trickier to implement while following a DFS exploration strategy. A key invariant of DFS exploration is that, during a concolic iteration, the previous path condition (used to generate the current input) is a prefix of the current path condition up to the negated constraint. This does not hold with rule C-DefaultError-Early, as the local constraint is dropped from the path condition. We leave the study of alternative exploration strategies and their impact as future work.
@@ 4.2 Generating Human-compatible Testcases
To do so, we rely on the encoding of optional constraints in the solver, dubbed “soft constraints” [59]. Z3 natively allows to combine SMT reasoning with solving optimization objectives, including the specification of soft constraints; however, we encountered significant slowdowns when attempting to use this feature, raising scalability issues, that have been reported to Z3 developers. We posit that this is due to non-linearity in our soft-constraints – e.g., when specifying – and in our rounding function defined in Sec. 3. We instead adopt an alternative approach, by implementing custom soft constraints through multiple queries to the solver. Concretely, when an iteration of concolic execution terminates and a new path condition is generated, we first query Z3 to check whether is satisfiable. If so, we further query Z3 by iteratively refining with additional constraints forcing human-friendlier input generation.
@@ 4.3 Pattern-Matching Case Folding
Consider for example the snippet in Fig. 11, extracted from the Catala implementation of the French housing benefits. At lines 5-13, this program performs a match on the 9 variants of a sum type representing geographic areas of France (mainland, or overseas territories). To analyze this program, a naive implementation of our concolic interpreter would therefore consider 9 different, disjoint branches, leading to 9 distinct iterations.
@@
[numbers=left,xleftmargin=2em,xrightmargin=0em,autogobble,firstline=2,lastline=15,firstnumber=1]catala_encode/pattern_matching_folding.catala_en
@@ 4.4 SMT-Solving Optimizations
Z3 provides incremental solving capabilities, where sets of constraints can be incrementally added or removed through stack-based push and pop operators. When popping the stack, the solver backtracks to an earlier proof state, only removing lemmas learned between push and pop and avoiding reproving many facts. We implemented support for incremental solving, pushing a new stack for each element in the path condition. As CUTECat’s exploration strategy is currently fixed to depth-first search (DFS), incremental solving is particularly appealing: constraints are naturally changed in a last-in-first-out fashion, corresponding to a stack-based data structure.